OpenStudy (anonymous):

Simplify (3x + 2)(4x - 5)

OpenStudy (anonymous):

If anyone is about to just give the answer, please read: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

(a+b)(c+d) = ac+ad+bc+bd

OpenStudy (anonymous):

Moreover: (a+b)*(c-d) = ac - ad + bc - bd

OpenStudy (anonymous):

so 12x -15 +6x - 10?

OpenStudy (anonymous):

look ,Multiply the first number (3x)to the second bracket =3x*(4x-5)=12x^2-15x. and then the second number to the same bracket=2*(4x-5)=8x-10 and then collect same terms =12x^2-15x+8x-10=12x^2-7x-10

OpenStudy (anonymous):

Almost right RiyaMoon, don't forget them x though! :)

OpenStudy (anonymous):

\[12x ^{2}-7x-10\]

OpenStudy (anonymous):

And if you have for example: 5x * 3x --> 15x^2 @jorge3o5 Read this: http://openstudy.com/code-of-conduct